OpenStudy (shamil98):

logarithmic differentiation how do i differentiate something like: y = 2x^e^x

OpenStudy (shamil98):

\[\large y= 2x^{e^x}\]

OpenStudy (anonymous):

take nature log for both sides and use implicit differrentication

OpenStudy (shamil98):

\[\ln y = e^x \ln 2x\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[ \large y= 2x^{e^x}\\ \ln y = \ln 2 + e^x \ln x \]

OpenStudy (shamil98):

then with respect to x, uh, and using the product rule.. \[\frac{ 1 }{ y }y' = \frac{ 1 }{ x } e^x + e^x \ln 2x\]

OpenStudy (anonymous):

Yes, now find y'

OpenStudy (shamil98):

\[y' = y(\frac{ e^x }{ x } + e^x \ln x)\]

OpenStudy (shamil98):

that should be a ln 2x*

OpenStudy (anonymous):

Replace y by its original formula

OpenStudy (shamil98):

\[\huge y' = 2x^{e^x}(\frac{ e^x }{ x } + e^x \ln 2x)\]